if the sum of two numbers is 2 and the difference is 12 what are the numbers?
Answer by sudhanshu_kmr(1152)


Let numbers are x and y.
then x+y = 2

and x-y = 12
now addition both equation we get 2x= 14
=> x = 14/2 = 7
then
y = 2 - x = 2 - 7 = -5

so, numbers are 7 and -5